\begin{thebibliography}{100}

\bibitem{begley2012drug}
Begley CG, Ellis LM.
\newblock Drug development: raise standards for preclinical cancer research.
\newblock Nature. 2012;483(7391):531--533.

\bibitem{prinz2011believe}
Prinz F, Schlange T, Asadullah K.
\newblock Believe it or not: how much can we rely on published data on
  potential drug targets?
\newblock Nature Reviews Drug Discovery. 2011;10(9):712--712.

\bibitem{freedman2015economics}
Freedman LP, Cockburn IM, Simcoe TS.
\newblock The economics of reproducibility in preclinical research.
\newblock PLoS Biol. 2015;13(6):e1002165.

\bibitem{chalmers2009avoidable}
Chalmers I, Glasziou P.
\newblock Avoidable waste in the production and reporting of research evidence.
\newblock Obstetrics \& Gynecology. 2009;114(6):1341--1345.

\bibitem{scott2008design}
Scott S, Kranz JE, Cole J, Lincecum JM, Thompson K, Kelly N, et~al.
\newblock Design, power, and interpretation of studies in the standard murine
  model of ALS.
\newblock Amyotrophic Lateral Sclerosis. 2008;9(1):4--15.

\bibitem{begley2013unappreciated}
Begley CG.
\newblock An unappreciated challenge to oncology drug discovery: pitfalls in
  preclinical research.
\newblock Am Soc Clin Oncol Educ Book. 2013;33:466--468.

\bibitem{steckler2015preclinical}
Steckler T.
\newblock Preclinical data reproducibility for R\&D -- the challenge for
  neuroscience.
\newblock Psychopharmacology. 2015;232(2):317--320.

\bibitem{de2015failure}
De~Los~Angeles A, Ferrari F, Fujiwara Y, Mathieu R, Lee S, Lee S, et~al.
\newblock Failure to replicate the STAP cell phenomenon.
\newblock Nature. 2015;525(7570):E6--E9.

\bibitem{kola2004can}
Kola I, Landis J.
\newblock Can the pharmaceutical industry reduce attrition rates?
\newblock Nature Reviews Drug Discovery. 2004;3(8):711--716.

\bibitem{macleod2014biomedical}
Macleod MR, Michie S, Roberts I, Dirnagl U, Chalmers I, Ioannidis JP, et~al.
\newblock Biomedical research: increasing value, reducing waste.
\newblock The Lancet. 2014;383(9912):101--104.

\bibitem{kyzas2007almost}
Kyzas PA, Denaxa-Kyza D, Ioannidis JP.
\newblock Almost all articles on cancer prognostic markers report statistically
  significant results.
\newblock European Journal of Cancer. 2007;43(17):2559--2579.

\bibitem{hirst2014need}
Hirst JA, Howick J, Aronson JK, Roberts N, Perera R, Koshiaris C, et~al.
\newblock The need for randomization in animal trials: an overview of
  systematic reviews.
\newblock PLoS ONE. 2014;9(6):e98856.

\bibitem{Hyman155cm11}
Hyman SE.
\newblock Revolution stalled.
\newblock Science Translational Medicine. 2012;4(155):155cm11.

\bibitem{miller2010pharma}
Miller G.
\newblock Is pharma running out of brainy ideas?
\newblock Science. 2010;329(5991):502--504.

\bibitem{schoenfeld2013everything}
Schoenfeld JD, Ioannidis JP.
\newblock Is everything we eat associated with cancer? A systematic cookbook
  review.
\newblock The American Journal of Clinical Nutrition. 2013;97(1):127--134.

\bibitem{ioannidis2005most}
Ioannidis JP.
\newblock Why most published research findings are false.
\newblock PLoS Med. 2005;2(8):e124.

\bibitem{ioannidis2013s}
Ioannidis J, Doucouliagos C.
\newblock What's to know about the credibility of empirical economics?
\newblock Journal of Economic Surveys. 2013;27(5):997--1004.

\bibitem{pankevich2014improving}
Pankevich DE, Altevogt BM, Dunlop J, Gage FH, Hyman SE.
\newblock Improving and accelerating drug development for nervous system
  disorders.
\newblock Neuron. 2014;84(3):546--553.

\bibitem{lowenstein2009uncertainty}
Lowenstein PR, Castro MG.
\newblock Uncertainty in the translation of preclinical experiments to clinical
  trials. Why do most phase III clinical trials fail?
\newblock Current Gene Therapy. 2009;9(5):368--374.

\bibitem{henderson2013threats}
Henderson VC, Kimmelman J, Fergusson D, Grimshaw JM, Hackam DG.
\newblock Threats to validity in the design and conduct of preclinical efficacy
  studies: a systematic review of guidelines for in vivo animal experiments.
\newblock PLoS Med. 2013;10(7):e1001489.

\bibitem{dirnagl2012international}
Dirnagl U, Fisher M.
\newblock International, multicenter randomized preclinical trials in
  translational stroke research: it's time to act.
\newblock Journal of Cerebral Blood Flow \& Metabolism. 2012;32(6):933--935.

\bibitem{dirnagl2009stroke}
Dirnagl U, Macleod MR.
\newblock Stroke research at a road block: the streets from adversity should be
  paved with meta-analysis and good laboratory practice.
\newblock British Journal of Pharmacology. 2009;157(7):1154--1156.

\bibitem{dirnagl2006bench}
Dirnagl U.
\newblock Bench to bedside: the quest for quality in experimental stroke
  research.
\newblock Journal of Cerebral Blood Flow \& Metabolism. 2006;26(12):1465--1478.

\bibitem{rosenblatt2016incentive}
Rosenblatt M.
\newblock An incentive-based approach for improving data reproducibility.
\newblock Science Translational Medicine. 2016;8(336):336ed5.

\bibitem{mobley2013survey}
Mobley A, Linder SK, Braeuer R, Ellis LM, Zwelling L.
\newblock A survey on data reproducibility in cancer research provides insights
  into our limited ability to translate findings from the laboratory to the
  clinic.
\newblock PLoS ONE. 2013;8(5):e63221.

\bibitem{open2015estimating}
{Open Science Collaboration}, et~al.
\newblock Estimating the reproducibility of psychological science.
\newblock Science. 2015;349(6251):aac4716.

\bibitem{ioannidis2005contradicted}
Ioannidis JP.
\newblock Contradicted and initially stronger effects in highly cited clinical
  research.
\newblock JAMA. 2005;294(2):218--228.

\bibitem{ioannidis2007non}
Ioannidis JP.
\newblock Non-replication and inconsistency in the genome-wide association
  setting.
\newblock Human Heredity. 2007;64(4):203--213.

\bibitem{steward2012replication}
Steward O, Popovich PG, Dietrich WD, Kleitman N.
\newblock Replication and reproducibility in spinal cord injury research.
\newblock Experimental Neurology. 2012;233(2):597--605.

\bibitem{chang2015economics}
Chang AC, Li P.
\newblock Is economics research replicable? Sixty published papers from
  thirteen journals say ``usually not''. 2015;\href
  {http://dx.doi.org/10.2139/ssrn.2669564} {doi:10.2139/ssrn.2669564}.

\bibitem{maxwell2015psychology}
Maxwell SE, Lau MY, Howard GS.
\newblock Is psychology suffering from a replication crisis? What does
  ``failure to replicate'' really mean?
\newblock American Psychologist. 2015;70(6):487.

\bibitem{anderson2016response}
Anderson CJ, Bahn{\'\i}k {\v{S}}, Barnett-Cowan M, Bosco FA, Chandler J,
  Chartier CR, et~al.
\newblock Response to comment on ``Estimating the reproducibility of
  psychological science''.
\newblock Science. 2016;351(6277):1037--1037.

\bibitem{stroebe2014alleged}
Stroebe W, Strack F.
\newblock The alleged crisis and the illusion of exact replication.
\newblock Perspectives on Psychological Science. 2014;9(1):59--71.

\bibitem{klein2014investigating}
Klein RA, Ratliff KA, Vianello M, Adams~Jr RB, Bahn{\'\i}k {\v{S}}, Bernstein
  MJ, et~al.
\newblock Investigating variation in replicability.
\newblock Social Psychology. 2014;45(3):142--152.

\bibitem{pashler2012replicability}
Pashler H, Harris CR.
\newblock Is the replicability crisis overblown? Three arguments examined.
\newblock Perspectives on Psychological Science. 2012;7(6):531--536.

\bibitem{camerer2016evaluating}
Camerer CF, Dreber A, Forsell E, Ho TH, Huber J, Johannesson M, et~al.
\newblock Evaluating replicability of laboratory experiments in economics.
\newblock Science. 2016;351(6280):1433--1436.

\bibitem{etz2016bayesian}
Etz A, Vandekerckhove J.
\newblock A Bayesian perspective on the reproducibility project: psychology.
\newblock PLoS ONE. 2016;11(2):e0149794.

\bibitem{michalek2010costs}
Michalek AM, Hutson AD, Wicher CP, Trump DL.
\newblock The costs and underappreciated consequences of research misconduct: a
  case study.
\newblock PLoS Med. 2010;7(8):e1000318.

\bibitem{fanelli2010positive}
Fanelli D.
\newblock ``Positive'' results increase down the hierarchy of the sciences.
\newblock PLoS ONE. 2010;5(4):e10068.

\bibitem{lilienfeld2012public}
Lilienfeld SO.
\newblock Public skepticism of psychology: why many people perceive the study
  of human behavior as unscientific.
\newblock American Psychologist. 2012;67(2):111.

\bibitem{cesario2014priming}
Cesario J.
\newblock Priming, replication, and the hardest science.
\newblock Perspectives on Psychological Science. 2014;9(1):40--48.

\bibitem{Pashler01112012}
Pashler H, Wagenmakers EJ.
\newblock Editors' introduction to the special section on replicability in
  psychological science: a crisis of confidence?
\newblock Perspectives on Psychological Science. 2012;7(6):528--530.

\bibitem{bakker2012rules}
Bakker M, van Dijk A, Wicherts JM.
\newblock The rules of the game called psychological science.
\newblock Perspectives on Psychological Science. 2012;7(6):543--554.

\bibitem{Bones01052012}
Bones AK.
\newblock We knew the future all along: scientific hypothesizing is much more
  accurate than other forms of precognition -- a satire in one part.
\newblock Perspectives on Psychological Science. 2012;7(3):307--309.

\bibitem{wagenmakers2011psychologists}
Wagenmakers EJ, Wetzels R, Borsboom D, van~der Maas HL.
\newblock Why psychologists must change the way they analyze their data: the
  case of psi: comment on Bem (2011).
\newblock Journal of Personality and Social Psychology. 2011;100(3):426--432.

\bibitem{ferguson2015everybody}
Ferguson CJ.
\newblock ``Everybody knows psychology is not a real science'': public
  perceptions of psychology and how we can improve our relationship with
  policymakers, the scientific community, and the general public.
\newblock American Psychologist. 2015;70(6):527.

\bibitem{goodman2016does}
Goodman SN, Fanelli D, Ioannidis JP.
\newblock What does research reproducibility mean?
\newblock Science Translational Medicine. 2016;8(341):341ps12.

\bibitem{brandt2014replication}
Brandt MJ, Ijzerman H, Dijksterhuis A, Farach FJ, Geller J, Giner-Sorolla R,
  et~al.
\newblock The replication recipe: what makes for a convincing replication?
\newblock Journal of Experimental Social Psychology. 2014;50:217--224.

\bibitem{clemens2015meaning}
Clemens MA.
\newblock The meaning of failed replications: a review and proposal.
\newblock Journal of Economic Surveys. 2015;\href
  {http://dx.doi.org/10.1111/joes.12139} {doi:10.1111/joes.12139}.

\bibitem{ioannidis2014make}
Ioannidis JP.
\newblock How to make more published research true.
\newblock PLoS Med. 2014;11(10):e1001747.

\bibitem{collins2014nih}
Collins FS, Tabak LA.
\newblock NIH plans to enhance reproducibility.
\newblock Nature. 2014;505(7485):612.

\bibitem{vesterinen2010improving}
Vesterinen HM, Sena ES, Williams A, Chandran S, Macleod MR, et~al.
\newblock Improving the translational hit of experimental treatments in
  multiple sclerosis.
\newblock Multiple Sclerosis. 2010;16(9):1044--1055.

\bibitem{asendorpf2013recommendations}
Asendorpf JB, Conner M, De~Fruyt F, De~Houwer J, Denissen JJ, Fiedler K, et~al.
\newblock Recommendations for increasing replicability in psychology.
\newblock European Journal of Personality. 2013;27(2):108--119.

\bibitem{nosek2015promoting}
Nosek BA, Alter G, Banks G, Borsboom D, Bowman S, Breckler S, et~al.
\newblock Promoting an open research culture.
\newblock Science. 2015;348(6242):1422--1425.

\bibitem{cumming2013new}
Cumming G.
\newblock The new statistics: why and how.
\newblock Psychological Science. 2014;25(1):7--29.

\bibitem{iqbal2016reproducible}
Iqbal SA, Wallach JD, Khoury MJ, Schully SD, Ioannidis JP.
\newblock Reproducible research practices and transparency across the
  biomedical literature.
\newblock PLoS Biol. 2016;14(1):e1002333.

\bibitem{holdren2013increasing}
Holdren JP.
\newblock Increasing access to the results of federally funded scientific
  research. Memorandum for the heads of executive departments and agencies.
\newblock Office of Science and Technology Policy, Executive Office of the
  President, Washington, DC. 2013;Available from:
  \url{https://www.whitehouse.gov/sites/default/files/microsites/ostp/ostp_public_access_memo_2013.pdf}.

\bibitem{miguel2014promoting}
Miguel E, Camerer C, Casey K, Cohen J, Esterling KM, Gerber A, et~al.
\newblock Promoting transparency in social science research.
\newblock Science. 2014;343(6166):30--31.

\bibitem{roche2014troubleshooting}
Roche DG, Lanfear R, Binning SA, Haff TM, Schwanz LE, Cain KE, et~al.
\newblock Troubleshooting public data archiving: suggestions to increase
  participation.
\newblock PLoS Biol. 2014;12(1):e1001779.

\bibitem{chalmers2014increase}
Chalmers I, Bracken MB, Djulbegovic B, Garattini S, Grant J, G{\"u}lmezoglu AM,
  et~al.
\newblock How to increase value and reduce waste when research priorities are
  set.
\newblock The Lancet. 2014;383(9912):156--165.

\bibitem{ioannidis2014increasing}
Ioannidis JP, Greenland S, Hlatky MA, Khoury MJ, Macleod MR, Moher D, et~al.
\newblock Increasing value and reducing waste in research design, conduct, and
  analysis.
\newblock The Lancet. 2014;383(9912):166--175.

\bibitem{salman2014increasing}
Salman RAS, Beller E, Kagan J, Hemminki E, Phillips RS, Savulescu J, et~al.
\newblock Increasing value and reducing waste in biomedical research regulation
  and management.
\newblock The Lancet. 2014;383(9912):176--185.

\bibitem{yordanov2015avoidable}
Yordanov Y, Dechartres A, Porcher R, Boutron I, Altman DG, Ravaud P.
\newblock Avoidable waste of research related to inadequate methods in clinical
  trials.
\newblock BMJ. 2015;350:h809.

\bibitem{moher2015four}
Moher D, Altman DG.
\newblock Four proposals to help improve the medical research literature.
\newblock PLoS Med. 2015;12(9):e1001864.

\bibitem{ioannidis2014assessing}
Ioannidis JP, Khoury MJ.
\newblock Assessing value in biomedical research: the PQRST of appraisal and
  reward.
\newblock JAMA. 2014;312(5):483--484.

\bibitem{vasilevsky2013reproducibility}
Vasilevsky NA, Brush MH, Paddock H, Ponting L, Tripathy SJ, LaRocca GM, et~al.
\newblock On the reproducibility of science: unique identification of research
  resources in the biomedical literature.
\newblock PeerJ. 2013;1:e148.

\bibitem{international2004international}
{International Committee of Medical Journal Editors}, et~al.
\newblock Uniform requirements for manuscripts submitted to biomedical
  journals: writing and editing for biomedical publication.
\newblock Haematologica. 2004;89(3):264.

\bibitem{kilkenny2010improving}
Kilkenny C, Browne WJ, Cuthill IC, Emerson M, Altman DG.
\newblock Improving bioscience research reporting: the ARRIVE guidelines for
  reporting animal research.
\newblock PLoS Biol. 2010;8(6):e1000412.

\bibitem{macleod2009reprint}
Macleod MR, Fisher M, O'Collins V, Sena ES, Dirnagl U, Bath PM, et~al.
\newblock Good laboratory practice: preventing introduction of bias at the
  bench.
\newblock Journal of Cerebral Blood Flow \& Metabolism. 2009;29(2):221--223.

\bibitem{tooth2005quality}
Tooth L, Ware R, Bain C, Purdie DM, Dobson A.
\newblock Quality of reporting of observational longitudinal research.
\newblock American Journal of Epidemiology. 2005;161(3):280--288.

\bibitem{festing2002guidelines}
Festing MF, Altman DG.
\newblock Guidelines for the design and statistical analysis of experiments
  using laboratory animals.
\newblock ILAR Journal. 2002;43(4):244--258.

\bibitem{moher2010guidance}
Moher D, Schulz KF, Simera I, Altman DG.
\newblock Guidance for developers of health research reporting guidelines.
\newblock PLoS Med. 2010;7(2):e1000217.

\bibitem{casadevall2012reforming}
Casadevall A, Fang FC.
\newblock Reforming science: methodological and cultural reforms.
\newblock Infection and Immunity. 2012;80(3):891--896.

\bibitem{nosek2012scientificI}
Nosek BA, Bar-Anan Y.
\newblock Scientific utopia: I. Opening scientific communication.
\newblock Psychological Inquiry. 2012;23(3):217--243.

\bibitem{nosek2012scientificII}
Nosek BA, Spies JR, Motyl M.
\newblock Scientific utopia: II. Restructuring incentives and practices to
  promote truth over publishability.
\newblock Perspectives on Psychological Science. 2012;7(6):615--631.

\bibitem{everett2015tragedy}
Everett JA, Earp BD.
\newblock A tragedy of the (academic) commons: interpreting the replication
  crisis in psychology as a social dilemma for early-career researchers.
\newblock Frontiers in Psychology. 2015;6.

\bibitem{nekrutenko2012next}
Nekrutenko A, Taylor J.
\newblock Next-generation sequencing data interpretation: enhancing
  reproducibility and accessibility.
\newblock Nature Reviews Genetics. 2012;13(9):667--672.

\bibitem{sandve2013ten}
Sandve GK, Nekrutenko A, Taylor J, Hovig E.
\newblock Ten simple rules for reproducible computational research.
\newblock PLoS Comput Biol. 2013;9(10):e1003285.

\bibitem{ioannidis2011improving}
Ioannidis JP, Khoury MJ.
\newblock Improving validation practices in ``omic'' research.
\newblock Science. 2011;334(6060):1230--1232.

\bibitem{pusztai2013reproducibility}
Pusztai L, Hatzis C, Andre F.
\newblock Reproducibility of research and preclinical validation: problems and
  solutions.
\newblock Nature Reviews Clinical Oncology. 2013;10(12):720--724.

\bibitem{valentine2011replication}
Valentine JC, Biglan A, Boruch RF, Castro FG, Collins LM, Flay BR, et~al.
\newblock Replication in prevention science.
\newblock Prevention Science. 2011;12(2):103--117.

\bibitem{kidwell2016badges}
Kidwell MC, Lazarevi{\'c} LB, Baranski E, Hardwicke TE, Piechowski S,
  Falkenberg LS, et~al.
\newblock Badges to acknowledge open practices: a simple, low-cost, effective
  method for increasing transparency.
\newblock PLoS Biol. 2016;14(5):e1002456.

\bibitem{koole2012rewarding}
Koole SL, Lakens D.
\newblock Rewarding replications: a sure and simple way to improve
  psychological science.
\newblock Perspectives on Psychological Science. 2012;7(6):608--614.

\bibitem{landis2012call}
Landis SC, Amara SG, Asadullah K, Austin CP, Blumenstein R, Bradley EW, et~al.
\newblock A call for transparent reporting to optimize the predictive value of
  preclinical research.
\newblock Nature. 2012;490(7419):187--191.

\bibitem{moher2010consort}
Moher D, Hopewell S, Schulz KF, Montori V, G{\o}tzsche PC, Devereaux P, et~al.
\newblock CONSORT 2010 explanation and elaboration: updated guidelines for
  reporting parallel group randomised trials.
\newblock Journal of Clinical Epidemiology. 2010;63(8):e1--e37.

\bibitem{moseley2014beyond}
Moseley ET, Hsu DJ, Stone DJ, Celi LA.
\newblock Beyond open big data: addressing unreliable research.
\newblock Journal of Medical Internet Research. 2014;16(11).

\bibitem{peers2012search}
Peers IS, Ceuppens PR, Harbron C.
\newblock In search of preclinical robustness.
\newblock Nature Reviews Drug Discovery. 2012;11(10).

\bibitem{peers2014can}
Peers IS, South MC, Ceuppens PR, Bright JD, Pilling E.
\newblock Can you trust your animal study data?
\newblock Nature Reviews Drug Discovery. 2014;13(7):560--560.

\bibitem{ram2013git}
Ram K.
\newblock Git can facilitate greater reproducibility and increased transparency
  in science.
\newblock Source Code for Biology and Medicine. 2013;8(1):1.

\bibitem{schooler2014metascience}
Schooler JW.
\newblock Metascience could rescue the ``replication crisis''.
\newblock Nature. 2014;515:9.

\bibitem{baker2014two}
Baker D, Lidster K, Sottomayor A, Amor S.
\newblock Two years later: journals are not yet enforcing the ARRIVE guidelines
  on reporting standards for pre-clinical animal studies.
\newblock PLoS Biol. 2014;12(1):e1001756.

\bibitem{vanpaemel2015we}
Vanpaemel W, Vermorgen M, Deriemaecker L, Storms G.
\newblock Are we wasting a good crisis? The availability of psychological
  research data after the storm.
\newblock Collabra. 2015;1(1).
\newblock \href {http://dx.doi.org/10.1525/collabra.13}
  {doi:10.1525/collabra.13}.

\bibitem{roche2015public}
Roche DG, Kruuk LE, Lanfear R, Binning SA.
\newblock Public data archiving in ecology and evolution: how well are we
  doing?
\newblock PLoS Biol. 2015;13(11):e1002295.

\bibitem{vines2014availability}
Vines TH, Albert AY, Andrew RL, D{\'e}barre F, Bock DG, Franklin MT, et~al.
\newblock The availability of research data declines rapidly with article age.
\newblock Current Biology. 2014;24(1):94--97.

\bibitem{wicherts2006poor}
Wicherts JM, Borsboom D, Kats J, Molenaar D.
\newblock The poor availability of psychological research data for reanalysis.
\newblock American Psychologist. 2006;61(7):726.

\bibitem{moher2016increasing}
Moher D, Glasziou P, Chalmers I, Nasser M, Bossuyt PM, Korevaar DA, et~al.
\newblock Increasing value and reducing waste in biomedical research: who's
  listening?
\newblock The Lancet. 2016;387(10027):1573--1586.

\bibitem{grant2013reporting}
Grant SP, Mayo-Wilson E, Melendez-Torres G, Montgomery P.
\newblock Reporting quality of social and psychological intervention trials: a
  systematic review of reporting guidelines and trial publications.
\newblock PLoS ONE. 2013;8(5):e65442.

\bibitem{joseph2013open}
Joseph H.
\newblock The open access movement grows up: taking stock of a revolution.
\newblock PLoS Biol. 2013;11(10):e1001686.

\bibitem{song2010dissemination}
Song F, Parekh S, Hooper L, Loke YK, Ryder J, Sutton AJ, et~al.
\newblock Dissemination and publication of research findings: an updated review
  of related biases.
\newblock Health Technol Assess. 2010;14(8):1--193.

\bibitem{prayle2012compliance}
Prayle AP, Hurley MN, Smyth AR.
\newblock Compliance with mandatory reporting of clinical trial results on
  ClinicalTrials.gov: cross sectional study.
\newblock BMJ. 2012;344:d7373.

\bibitem{clarke2010clinical}
Clarke M, Hopewell S, Chalmers I.
\newblock Clinical trials should begin and end with systematic reviews of
  relevant evidence: 12 years and waiting.
\newblock The Lancet. 2010;376(9734):20--21.

\bibitem{florez2016bias}
Fl{\'o}rez-Vargas O, Brass A, Karystianis G, Bramhall M, Stevens R, Cruickshank
  S, et~al.
\newblock Bias in the reporting of sex and age in biomedical research on mouse
  models.
\newblock ELife. 2016;5:e13615.

\bibitem{bramhall2015quality}
Bramhall M, Fl{\'o}rez-Vargas O, Stevens R, Brass A, Cruickshank S.
\newblock Quality of methods reporting in animal models of colitis.
\newblock Inflammatory Bowel Diseases. 2015;21(6):1248--1259.

\bibitem{plint2006does}
Plint AC, Moher D, Morrison A, Schulz K, Altman DG, Hill C, et~al.
\newblock Does the CONSORT checklist improve the quality of reports of
  randomised controlled trials? A systematic review.
\newblock Medical Journal of Australia. 2006;185(5):263.

\bibitem{begley2015reproducibility}
Begley CG, Ioannidis JP.
\newblock Reproducibility in science: improving the standard for basic and
  preclinical research.
\newblock Circulation Research. 2015;116(1):116--126.

\bibitem{simmons2011false}
Simmons JP, Nelson LD, Simonsohn U.
\newblock False-positive psychology: undisclosed flexibility in data collection
  and analysis allows presenting anything as significant.
\newblock Psychological Science. 2011;\href
  {http://dx.doi.org/10.1177/0956797611417632} {doi:10.1177/0956797611417632}.

\bibitem{fiedler2011voodoo}
Fiedler K.
\newblock Voodoo correlations are everywhere -- not only in neuroscience.
\newblock Perspectives on Psychological Science. 2011;6(2):163--171.

\bibitem{ferguson2012vast}
Ferguson CJ, Heene M.
\newblock A vast graveyard of undead theories: publication bias and
  psychological science's aversion to the null.
\newblock Perspectives on Psychological Science. 2012;7(6):555--561.

\bibitem{young2008current}
Young NS, Ioannidis JP, Al-Ubaydli O.
\newblock Why current publication practices may distort science.
\newblock PLoS Med. 2008;5(10):e201.

\bibitem{tsilidis2013evaluation}
Tsilidis KK, Panagiotou OA, Sena ES, Aretouli E, Evangelou E, Howells DW,
  et~al.
\newblock Evaluation of excess significance bias in animal studies of
  neurological diseases.
\newblock PLoS Biol. 2013;11(7):e1001609.

\bibitem{brembs2013deep}
Brembs B, Munaf{\`o} M.
\newblock Deep impact: unintended consequences of journal rank. 2013;\href
  {http://arxiv.org/abs/1301.3748} {arXiv:1301.3748}.

\bibitem{steen2013has}
Steen RG, Casadevall A, Fang FC.
\newblock Why has the number of scientific retractions increased?
\newblock PLoS ONE. 2013;8(7):e68397.

\bibitem{oboyle2014chrysalis}
O'Boyle EH, Banks GC, Gonzalez-Mul{\'e} E.
\newblock The Chrysalis Effect: how ugly initial results metamorphosize into
  beautiful articles.
\newblock Journal of Management. 2014;\href
  {http://dx.doi.org/10.1177/0149206314527133} {doi:10.1177/0149206314527133}.

\bibitem{stamatakis2013undue}
Stamatakis E, Weiler R, Ioannidis J.
\newblock Undue industry influences that distort healthcare research, strategy,
  expenditure and practice: a review.
\newblock European Journal of Clinical Investigation. 2013;43(5):469--475.

\bibitem{sena2010publication}
Sena ES, Van Der~Worp HB, Bath PM, Howells DW, Macleod MR.
\newblock Publication bias in reports of animal stroke studies leads to major
  overstatement of efficacy.
\newblock PLoS Biol. 2010;8(3):e1000344.

\bibitem{mathieu2009comparison}
Mathieu S, Boutron I, Moher D, Altman DG, Ravaud P.
\newblock Comparison of registered and published primary outcomes in randomized
  controlled trials.
\newblock JAMA. 2009;302(9):977--984.

\bibitem{hannink2013comparison}
Hannink G, Gooszen HG, Rovers MM.
\newblock Comparison of registered and published primary outcomes in randomized
  clinical trials of surgical interventions.
\newblock Annals of Surgery. 2013;257(5):818--823.

\bibitem{crowe2015patients}
Crowe S, Fenton M, Hall M, Cowan K, Chalmers I.
\newblock Patients', clinicians' and the research communities' priorities for
  treatment research: there is an important mismatch.
\newblock Research Involvement and Engagement. 2015;1(1):1--10.
\newblock \href {http://dx.doi.org/10.1186/s40900-015-0003-x}
  {doi:10.1186/s40900-015-0003-x}.

\bibitem{blumle2016fate}
Bl{\"u}mle A, Meerpohl JJ, Schumacher M, von Elm E.
\newblock Fate of clinical research studies after ethical approval -- follow-up
  of study protocols until publication.
\newblock In: Ethics and Governance of Biomedical Research. Springer; 2016. p.
  109--123.

\bibitem{chan2014increasing}
Chan AW, Song F, Vickers A, Jefferson T, Dickersin K, G{\o}tzsche PC, et~al.
\newblock Increasing value and reducing waste: addressing inaccessible
  research.
\newblock The Lancet. 2014;383(9913):257--266.

\bibitem{glasziou2014reducing}
Glasziou P, Altman DG, Bossuyt P, Boutron I, Clarke M, Julious S, et~al.
\newblock Reducing waste from incomplete or unusable reports of biomedical
  research.
\newblock The Lancet. 2014;383(9913):267--276.

\bibitem{glasziou2014role}
Glasziou P.
\newblock The role of open access in reducing waste in medical research.
\newblock PLoS Med. 2014;11(5):e1001651.

\bibitem{hoffmann2012scatter}
Hoffmann T, Erueti C, Thorning S, Glasziou P.
\newblock The scatter of research: cross sectional comparison of randomised
  trials and systematic reviews across specialties.
\newblock BMJ. 2012;344(3223).

\bibitem{franco2014publication}
Franco A, Malhotra N, Simonovits G.
\newblock Publication bias in the social sciences: unlocking the file drawer.
\newblock Science. 2014;345(6203):1502--1505.

\bibitem{button2013power}
Button KS, Ioannidis JP, Mokrysz C, Nosek BA, Flint J, Robinson ES, et~al.
\newblock Power failure: why small sample size undermines the reliability of
  neuroscience.
\newblock Nature Reviews Neuroscience. 2013;14(5):365--376.

\bibitem{dwan2008systematic}
Dwan K, Altman DG, Arnaiz JA, Bloom J, Chan AW, Cronin E, et~al.
\newblock Systematic review of the empirical evidence of study publication bias
  and outcome reporting bias.
\newblock PLoS ONE. 2008;3(8):e3081.

\bibitem{ioannidis2008most}
Ioannidis JP.
\newblock Why most discovered true associations are inflated.
\newblock Epidemiology. 2008;19(5):640--648.

\bibitem{van2010can}
Van~der Worp HB, Howells DW, Sena ES, Porritt MJ, Rewell S, O'Collins V, et~al.
\newblock Can animal models of disease reliably inform human studies?
\newblock PLoS Med. 2010;7(3):e1000245.

\bibitem{ioannidis2012science}
Ioannidis JP.
\newblock Why science is not necessarily self-correcting.
\newblock Perspectives on Psychological Science. 2012;7(6):645--654.

\bibitem{stroebe2012scientific}
Stroebe W, Postmes T, Spears R.
\newblock Scientific misconduct and the myth of self-correction in science.
\newblock Perspectives on Psychological Science. 2012;7(6):670--688.

\bibitem{wagenmakers2012agenda}
Wagenmakers EJ, Wetzels R, Borsboom D, van~der Maas HL, Kievit RA.
\newblock An agenda for purely confirmatory research.
\newblock Perspectives on Psychological Science. 2012;7(6):632--638.

\bibitem{john2012measuring}
John LK, Loewenstein G, Prelec D.
\newblock Measuring the prevalence of questionable research practices with
  incentives for truth telling.
\newblock Psychological Science. 2012;\href
  {http://dx.doi.org/10.1177/0956797611430953} {doi:10.1177/0956797611430953}.

\bibitem{fanelli2011negative}
Fanelli D.
\newblock Negative results are disappearing from most disciplines and
  countries.
\newblock Scientometrics. 2011;90(3):891--904.

\bibitem{alberts2014rescuing}
Alberts B, Kirschner MW, Tilghman S, Varmus H.
\newblock Rescuing US biomedical research from its systemic flaws.
\newblock Proceedings of the National Academy of Sciences.
  2014;111(16):5773--5777.

\bibitem{ioannidis2014publication}
Ioannidis JP, Munafo MR, Fusar-Poli P, Nosek BA, David SP.
\newblock Publication and other reporting biases in cognitive sciences:
  detection, prevalence, and prevention.
\newblock Trends in Cognitive Sciences. 2014;18(5):235--241.

\bibitem{chambers2014instead}
Chambers CD, Feredoes E, Muthukumaraswamy SD, Etchells P.
\newblock Instead of ``playing the game'' it is time to change the rules:
  registered reports at AIMS Neuroscience and beyond.
\newblock AIMS Neuroscience. 2014;1(1):4--17.

\bibitem{glasziou2005paths}
Glasziou P, Haynes B.
\newblock The paths from research to improved health outcomes.
\newblock Evidence Based Nursing. 2005;8(2):36--38.

\bibitem{duff2010adequacy}
Duff JM, Leather H, Walden EO, LaPlant KD, George TJ.
\newblock Adequacy of published oncology randomized controlled trials to
  provide therapeutic details needed for clinical application.
\newblock Journal of the National Cancer Institute. 2010;102(10):702--705.

\bibitem{dancey2010quality}
Dancey JE.
\newblock From quality of publication to quality of care: translating trials to
  practice.
\newblock Journal of the National Cancer Institute. 2010;102(10):670--671.

\bibitem{kilkenny2009survey}
Kilkenny C, Parsons N, Kadyszewski E, Festing MF, Cuthill IC, Fry D, et~al.
\newblock Survey of the quality of experimental design, statistical analysis
  and reporting of research using animals.
\newblock PLoS ONE. 2009;4(11):e7824.

\bibitem{mcglynn2003quality}
McGlynn EA, Asch SM, Adams J, Keesey J, Hicks J, DeCristofaro A, et~al.
\newblock The quality of health care delivered to adults in the United States.
\newblock New England Journal of Medicine. 2003;348(26):2635--2645.

\bibitem{lemon2005surveying}
Lemon R, Dunnett SB.
\newblock Surveying the literature from animal experiments.
\newblock BMJ. 2005;330(7498):977--978.

\bibitem{ioannidis2001completeness}
Ioannidis JP, Lau J.
\newblock Completeness of safety reporting in randomized trials: an evaluation
  of 7 medical areas.
\newblock JAMA. 2001;285(4):437--443.

\bibitem{savovic2012influence}
Savovi{\'c} J, Jones HE, Altman DG, Harris RJ, J{\"u}ni P, Pildal J, et~al.
\newblock Influence of reported study design characteristics on intervention
  effect estimates from randomized, controlled trials.
\newblock Annals of Internal Medicine. 2012;157(6):429--438.

\bibitem{turner2008selective}
Turner EH, Matthews AM, Linardatos E, Tell RA, Rosenthal R.
\newblock Selective publication of antidepressant trials and its influence on
  apparent efficacy.
\newblock New England Journal of Medicine. 2008;358(3):252--260.

\bibitem{bero1998closing}
Bero LA, Grilli R, Grimshaw JM, Harvey E, Oxman AD, Thomson MA.
\newblock Closing the gap between research and practice: an overview of
  systematic reviews of interventions to promote the implementation of research
  findings.
\newblock BMJ. 1998;317(7156):465--469.

\bibitem{ramagopalan2014prevalence}
Ramagopalan S, Skingsley AP, Handunnetthi L, Klingel M, Magnus D, Pakpoor J,
  et~al.
\newblock Prevalence of primary outcome changes in clinical trials registered
  on ClinicalTrials.gov: a cross-sectional study.
\newblock F1000Research. 2014;3(77).
\newblock \href {http://dx.doi.org/10.12688/f1000research.3784.1}
  {doi:10.12688/f1000research.3784.1}.

\bibitem{williamson2012driving}
Williamson P, Altman D, Blazeby J, Clarke M, Gargon E.
\newblock Driving up the quality and relevance of research through the use of
  agreed core outcomes.
\newblock Journal of Health Services Research \& Policy. 2012;17(1):1--2.

\bibitem{stephan1996economics}
Stephan PE.
\newblock The economics of science.
\newblock Journal of Economic Literature. 1996;34(3):1199--1235.

\bibitem{knuteson2016knowledge}
Knuteson B.
\newblock Blunt honesty, incentives, and knowledge exchange. 2016;\href
  {http://arxiv.org/abs/1602.06188} {arXiv:1602.06188}.

\bibitem{lykken1968statistical}
Lykken DT.
\newblock Statistical significance in psychological research.
\newblock Psychological Bulletin. 1968;70(3p1):151.

\bibitem{elms1975crisis}
Elms AC.
\newblock The crisis of confidence in social psychology.
\newblock American Psychologist. 1975;30(10):967.

\bibitem{greenwald1975consequences}
Greenwald AG.
\newblock Consequences of prejudice against the null hypothesis.
\newblock Psychological Bulletin. 1975;82(1):1.

\bibitem{rosenthal1979file}
Rosenthal R.
\newblock The file drawer problem and tolerance for null results.
\newblock Psychological Bulletin. 1979;86(3):638.

\bibitem{altman1994scandal}
Altman DG.
\newblock The scandal of poor medical research.
\newblock BMJ. 1994;308(6924):283--284.

\bibitem{hackam2006translation}
Hackam DG, Redelmeier DA.
\newblock Translation of research evidence from animals to humans.
\newblock JAMA. 2006;296(14):1727--1732.

\bibitem{pocock1987statistical}
Pocock SJ, Hughes MD, Lee RJ.
\newblock Statistical problems in the reporting of clinical trials.
\newblock New England Journal of Medicine. 1987;317(7):426--432.

\bibitem{sterling1995publication}
Sterling TD, Rosenbaum WL, Weinkam JJ.
\newblock Publication decisions revisited: the effect of the outcome of
  statistical tests on the decision to publish and vice versa.
\newblock The American Statistician. 1995;49(1):108--112.

\bibitem{vul2009puzzlingly}
Vul E, Harris C, Winkielman P, Pashler H.
\newblock Puzzlingly high correlations in fMRI studies of emotion, personality,
  and social cognition.
\newblock Perspectives on Psychological Science. 2009;4(3):274--290.

\bibitem{easterbrook1991publication}
Easterbrook PJ, Gopalan R, Berlin J, Matthews DR.
\newblock Publication bias in clinical research.
\newblock The Lancet. 1991;337(8746):867--872.

\bibitem{kerr1998harking}
Kerr NL.
\newblock HARKing: hypothesizing after the results are known.
\newblock Personality and Social Psychology Review. 1998;2(3):196--217.

\bibitem{knuteson2011capitalist}
Knuteson B.
\newblock Capitalist science. 2011;\href {http://arxiv.org/abs/1102.2474}
  {arXiv:1102.2474}.

\end{thebibliography}

\documentclass[twocolumn,twoside,prd,floatfix,letterpaper]{revtex4}

\usepackage{ifpdf}
\ifpdf 
  \usepackage[pdftex]{graphicx} 
\else 
  \usepackage{graphicx} 
\fi
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{fancyhdr}
\usepackage{color}
\usepackage{rotating}
\usepackage{natbib}
\usepackage[colorlinks,hyperindex]{hyperref}
\usepackage{breakurl}

\begin{document} 

\title{The Solution to Science's Replication Crisis}
\author{Bruce Knuteson}
\homepage{http://bruceknuteson.com}
\email{knuteson@kn-x.com}

\begin{abstract}
The solution to science's replication crisis is a new ecosystem in which scientists sell what they learn from their research.  In each pairwise transaction, the information seller makes (loses) money if he turns out to be correct (incorrect).  Responsibility for the determination of correctness is delegated, with appropriate incentives, to the information purchaser.  Each transaction is brokered by a central exchange, which holds money from the anonymous information buyer and anonymous information seller in escrow, and which enforces a set of incentives facilitating the transfer of useful, bluntly honest information from the seller to the buyer.  This new ecosystem, {\em{capitalist science}}, directly addresses socialist science's replication crisis by explicitly rewarding accuracy and penalizing inaccuracy.  

The problematic incentives within socialist science that have led to socialist science's replication crisis are too embedded to remove.  Those seeking accurate knowledge to inform important, high stakes decisions will gradually start to rely on capitalist science, which penalizes the selling of information that turns out to be wrong.  Researchers with genuinely valuable information will choose to sell it rather than publish it away for free.  The most useful, interesting, and accurate information will therefore gradually move from the ecosystem of socialist science to the ecosystem of capitalist science, where it is properly valued.  

Capitalist science is the solution to socialist science's replication crisis.
\end{abstract}

\maketitle

\section{Socialist Science}
\label{sec:SocialistScience}

Drug company Q reads a scholarly article in a distinguished journal authored by academic A.  Motivated in significant part by A's findings, Q starts a new research effort in this seemingly promising direction.

One year and one million dollars later, Q realizes A's findings are basically wrong~\cite{begley2012drug,prinz2011believe,freedman2015economics,chalmers2009avoidable,scott2008design,begley2013unappreciated,steckler2015preclinical,de2015failure,kola2004can,macleod2014biomedical,kyzas2007almost,hirst2014need,Hyman155cm11,miller2010pharma,schoenfeld2013everything}.

Variants of this story may be invoked by a convenient buzz phrase:  science's ``replication crisis,'' ``replicability crisis,'' or ``reproducibility crisis.''  Q is not always a drug company.  A is usually an academic.

Q, the sucker in this story, realizes he is continually getting shafted~\cite{ioannidis2005most,ioannidis2013s,pankevich2014improving,lowenstein2009uncertainty,henderson2013threats,dirnagl2012international,dirnagl2009stroke,dirnagl2006bench,rosenblatt2016incentive}.  Systematic attempts to repeat many published studies fail~\cite{mobley2013survey,open2015estimating,ioannidis2005contradicted,ioannidis2007non,steward2012replication,chang2015economics}, spreading awareness to the masses and stoking discontent.  Debates erupt over the scale of the problem~\cite{maxwell2015psychology,anderson2016response,stroebe2014alleged,klein2014investigating,pashler2012replicability,camerer2016evaluating,etz2016bayesian,michalek2010costs}.  Some hope extermination of psychology departments will be sufficient~\cite{fanelli2010positive,lilienfeld2012public,cesario2014priming,Pashler01112012,bakker2012rules,Bones01052012,wagenmakers2011psychologists,ferguson2015everybody}.  Others work to re(de)fine the notion of ``replication''~\cite{goodman2016does,brandt2014replication,clemens2015meaning}.  Incremental improvements to the current system are proposed~\cite{ioannidis2014make,collins2014nih,vesterinen2010improving,asendorpf2013recommendations,nosek2015promoting,cumming2013new,iqbal2016reproducible,holdren2013increasing,miguel2014promoting,roche2014troubleshooting,chalmers2014increase,ioannidis2014increasing,salman2014increasing,yordanov2015avoidable,moher2015four,ioannidis2014assessing,vasilevsky2013reproducibility,international2004international,kilkenny2010improving,macleod2009reprint,tooth2005quality,festing2002guidelines,moher2010guidance,casadevall2012reforming,nosek2012scientificI,nosek2012scientificII,everett2015tragedy,nekrutenko2012next,sandve2013ten,ioannidis2011improving,pusztai2013reproducibility,valentine2011replication,kidwell2016badges,koole2012rewarding,landis2012call,moher2010consort,moseley2014beyond,peers2012search,peers2014can,ram2013git,schooler2014metascience}~\footnote{{\em{Open science}}, or {\em{even-more-socialist science}}, is a common theme among proposed incremental improvements.  The idea of encouraging more transparent access to data and analysis code is an attractive one.  We ourselves pushed it in particle physics -- hard, and for many years.  Unfortunately, if unsurprisingly, incentives are simply too misaligned for it to work.  If we, as a society, want more of something -- like apples, say, or knowledge about how nature works -- we may be better off making it easy for people who produce apples to sell them than mandating that they make their orchards available for anyone to pick.}.  Most are never enacted~\cite{baker2014two}.  The rest have limited impact~\cite{vanpaemel2015we,roche2015public,vines2014availability,wicherts2006poor,moher2016increasing,grant2013reporting,joseph2013open,song2010dissemination,prayle2012compliance,clarke2010clinical,florez2016bias,bramhall2015quality,plint2006does}.

The incentive structures contributing to the replication crisis are observed to be deeply entrenched features of the ecosystem within which science is done~\cite{begley2015reproducibility,simmons2011false,fiedler2011voodoo,ferguson2012vast,young2008current,tsilidis2013evaluation,brembs2013deep,steen2013has,oboyle2014chrysalis,stamatakis2013undue,sena2010publication,mathieu2009comparison,hannink2013comparison,crowe2015patients,blumle2016fate,chan2014increasing,glasziou2014reducing,glasziou2014role,hoffmann2012scatter,franco2014publication,button2013power,dwan2008systematic,ioannidis2008most,van2010can,ioannidis2012science,stroebe2012scientific,wagenmakers2012agenda,john2012measuring,fanelli2011negative,alberts2014rescuing,ioannidis2014publication,chambers2014instead} and subsequently applied~\cite{glasziou2005paths,duff2010adequacy,dancey2010quality,kilkenny2009survey,mcglynn2003quality,lemon2005surveying,ioannidis2001completeness,savovic2012influence,turner2008selective,bero1998closing,ramagopalan2014prevalence,williamson2012driving}.  This ecosystem comprises institutions (including funding agencies, industry, universities, and academic journals), culture, accepted practice (including procedures by which grant money is awarded, articles published, promotions granted, and accolades bestowed), and the bureaucracy required to support all of these.  The ecosystem is complex, multifaceted, and not easily changed.  The incentive structures contributing to the replication crisis are, similarly, not easily changed.  A timely solution to the replication crisis probably requires a new ecosystem.

For convenience, we refer to the current ecosystem~\cite{stephan1996economics} -- funded by taxpayers, with published results available to all citizens -- as {\em{socialist science}}~\footnote{The phrase ``socialist science'' is obviously a grotesquely crude caricature of the intricate and often nuanced set of incentives joining the actors in the current science ecosystem.  We intend the phrase as a neutral description of an aspect of the current ecosystem germane to the present discussion.  A reader who dislikes the phrase is encouraged to mentally replace it with ``the ecosystem within which science is currently carried out,'' or some alternative reference thereto.}~\footnote{We focus on the incentive flaws common to all socialist science, ignoring differences among scientific disciplines.  Rather than treat the symptoms, which express differently in the social sciences, life sciences, and physical sciences, we focus on the underlying disease.}.

\section{Capitalist Science}
\label{sec:CapitalistScience}

Perhaps we can construct a different ecosystem.  

Lacking imagination, we look to see what other professions do.  Many professions exhibit a very interesting behavior.  There occur {\em{pairwise transactions}}, in which one party {\em{sells}} something of value to another party.  One party (A) gives something of value to some other party (Q).  In return, Q gives A something called {\em{money}}.

Perhaps our academic (A) can sell what he learns to those (Q) who find A's information valuable.  Many deep pocketed industries, including the pharmaceutical industry, make high stakes decisions motivated by academic research.  A should have little difficulty finding customers Q.

To incentivize accuracy, we need some sort of transaction that causes A to lose money if he turns out to be wrong.  We therefore need transactions for which the (in)accuracy of A's information can eventually be objectively determined, and we need a reliable, low-cost procedure for making that determination.

Suppose Q has a specific question and is willing to pay for a useful answer.  Q can ensure the usefulness of any answer he receives by specifying, explicitly or categorically, the set of allowed possible answers.  A third party (X) brokers the transaction, holds money from Q and A in escrow, and acts as arbiter.  If A provides an answer outside Q's set of allowed possible answers, A loses money.

In many cases of practical interest, Q can be tasked with determining the accuracy of A's answer.  If Q wants to know what chemical matter binds to a particular target and A answers with a specific molecule, Q will eventually know whether A's answer was right or wrong, and Q can provide evidence to this effect to arbiter X.  X collects a deposit from Q at the beginning of the transaction, and X returns this deposit to Q when Q provides sufficient evidence by a previously agreed upon date.  Q pays the same amount whether A's answer turns out to be right or wrong.  X's fee for brokering the transaction is independent of whether A's answer turns out to be right or wrong and whether or not Q provides evidence X considers sufficient.  A has an explicit monetary incentive to be accurate and an explicit monetary disincentive to provide information unless he is pretty sure he is correct.  Q has an explicit monetary incentive to ask a question that will be directly useful to him, and for which he can eventually determine, with reasonable proof, the accuracy of whatever answer A provides.  X, as an ongoing business interest, has a clear monetary incentive to thoughtfully serve its role as a reasonable and objective adjudicator of the evidence Q provides.  Q, A, and X all have an interest in ensuring Q's question is carefully specified to avoid subsequent ambiguities as to A's correctness.  This transaction protocol enforces a set of incentives facilitating paid, arm's length transfer of useful, bluntly honest information from A to Q.

Using this transaction protocol, introduced for the first time in Ref.~\cite{knuteson2016knowledge} and implemented at Ref.~\footnote{Kn-X; \url{http://kn-x.com}.  Patent pending.}, scientists can sell what they learn from their research.  

For convenience, we refer to this new ecosystem as {\em{capitalist science}}~\footnote{The phrase ``capitalist science'' is intended as a neutral description of a salient feature of this new ecosystem.  A reader who dislikes the phrase is encouraged to mentally replace it with ``the new ecosystem,'' or something similar.}~\footnote{The ``capitalist science'' in this article supersedes that of Ref.~\cite{knuteson2011capitalist}, which in retrospect is more of a hybrid between socialist and capitalist science.}~\footnote{To be clear, our use of the term ``capitalist'' does not arise from a belief, seemingly held by many, that free markets are the optimal solution to all problems, nor from a belief that global financial markets at the time of this writing function well and should be emulated.  (You have no idea.)  The complexity and frequent opacity of today's capitalism highlights the glaring need for a mechanism facilitating useful, bluntly honest information transfer between remote parties.   Given the embarrassing, hidden-in-plain-sight, farcically tragic comedy of errors that is recent financial history, a mechanism facilitating useful, bluntly honest, arm's length information transfer may turn out to be our best shot at saving capitalism \ldots\ or at least at sending a few of those responsible to prison next time around.}.

In capitalist science, A makes money only if he turns out to be correct.  A loses money if he turns out to be wrong.  Every transaction includes a monetary incentive for Q to determine the accuracy of A's answer and to back this determination with evidence deemed sufficient by an objective third party (X).  These features directly address socialist science's reproducibility crisis.  These incentives, present in capitalist science, explicitly and directly reward accuracy and penalize inaccuracy.  The absence of such explicit and direct incentives in socialist science is the reason socialist science is suffering a reproducibility crisis.

Capitalist science is the solution to socialist science's reproducibility crisis.

\section{Summary}
\label{sec:Summary}

Now that the solution to socialist science's reproducibility crisis has been found, we can roll up our sleeves \ldots\ sit back, relax, and let events unfold.

Socialist science will continue much as it has, producing results of similar quality.  Serious, thoughtful, and impassioned attempts will be made to save it.  Some of these may appear to work for a while, but ultimately they will fail.  The reproducibility crisis has been stewing for decades~\cite{lykken1968statistical,elms1975crisis,greenwald1975consequences,rosenthal1979file,altman1994scandal,hackam2006translation,pocock1987statistical,sterling1995publication,vul2009puzzlingly,easterbrook1991publication,kerr1998harking}.  The relevant malincentives are too firmly embedded in its large and unwieldy ecosystem for socialist science to produce a meaningful solution from within.

The key to change is Q, the socialist science sucker from Section~\ref{sec:SocialistScience}.  Q is smart.  Q is a socialist science sucker only because socialist science fails to sufficiently discourage the publication of inaccurate information by failing to hold A sufficiently accountable for being wrong.  Q is a socialist science sucker only because Q has had no other option.

Q now has another option.  Before launching a one year, one million dollar research project motivated by a result produced by socialist science, Q can spend ten days and ten thousand dollars on a few questions to check it out.  Before investing ten million dollars on a new materials science innovation, venture capitalists can anonymously conduct due diligence with a convenience and at a depth hitherto unimaginable.  Relying on free information of questionable accuracy produced by socialist science can be very costly.  Q is smart; Q is accustomed to paying for things of value; and Q's stakes are high.

Over time, Q increasingly relies on information obtained from capitalist science, where A has skin in the game.  When A has truly valuable information, he finds himself more interested in selling it than publishing it.  For less robust results, A is more inclined to publish, avoiding the harsh penalty imposed by capitalist science for providing information that turns out to be wrong.  Over time, useful, robust information slowly migrates to capitalist science, where it is appropriately valued.  Socialist science continues to publish the results A is willing to give away for free.

The forces in capitalist science's favor are so strong that socialist science should be allowed to fade into irrelevance in its natural course~\footnote{Although the information market unleashed by capitalist science could create millions of new science-related jobs, it would be irresponsible to reduce funding to socialist science until that promise has been realized.}.  Socialist science has been a remarkable institution.  The knowledge it has provided has been extraordinary in power and in scope.  Its use of an incentive structure spurned by the most developed of today's modern economies only makes its achievements all the more impressive.

We fully hope and expect socialist science will linger among us for many years.  At some deep level, the purpose and goals of socialist science are fundamentally different from those of capitalist science.  Capitalist science is designed to facilitate useful, accurate information transfer, and to provide A with a strong incentive to learn information some Q will find valuable.  Socialist science, in contrast, is not.

\bibliography{crisis_solution}

\end{document}